Hello, I am solving some questions in my textbook and I have seem to fallen into a slump. Please help.
A bacteria called escherichia coli are commonly found in the bladder. Suppose that 3000 of the bacteria are present at time t=0. Then t minutes later, the number of bacteria present can be approximated by n(t)=3000(2)^(t/20) (the t/20 is written in exponent form)
How many bacteria will be present after 10min? 20min? 30min? 40min? 60min?
If I could have someone workout the first few and I can probably get the rest done.

Answer by stanbon(57203)

Your equation says:
#of bacteria at time "t" is n(t) = 3000(2)^(t/20)
By the way this means the number doubles every 20 minutes.
n(10)=3000 (2)^(10/20)
n(10)=3000 (2)^(1/2)
n(10)=3000 (sqrt 2)
n(10)=4242.64....
So you see you have to substitute 20 or 30 or 40 or 60 into
the formula in place of the "n".
Cheers,
Stan H.